OpenStudy (anonymous):

What is the probability of getting a 6 each time if a die is rolled 4 times?

OpenStudy (kirbykirby):

\[\frac{1}{6}*\frac{1}{6}*\frac{1}{6}*\frac{1}{6}\]

OpenStudy (kirbykirby):

each roll of getting a "6" is just 1/6. Since you want a 6 on every roll, you want prob(die=6) AND prob(die=6) AND prob(die=6) AND prob(die=6)... so the "AND" tells you to use the product rule

OpenStudy (anonymous):

what is the product rule

OpenStudy (kirbykirby):

Product rule: Suppose you can do job 1 in p ways, and job 2 in q ways. Then, we can do job 1 AND job 2 in \(p*q\) ways Addition rule: Suppose we can do job 1 in p ways and job 2 in q ways. Then we can do either job 1 OR job 2 in \(p+q\) ways

OpenStudy (anonymous):

oh okay thank you